OpenStudy (anonymous):

f(x) = x2 – 9, g(x) = 2x + 3 and h(x) = x - 3. Find each of the following. a) (f + g)(x) b) (f/h)(x) c) g(f(x)) d) f(g(h(x))) e) (g – h)(8) f) h(f(-2))

OpenStudy (anonymous):

all at once?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1) put a plus sign between them, combine like terms

OpenStudy (anonymous):

2) put one on top of the other using a fraction bar

OpenStudy (anonymous):

a)x^2 +2x-6. i think thats the answer to a)

OpenStudy (anonymous):

c) requires some work \[g(f(x))=g(x^2-9)=2(x^2-9)+3\] then distribute and combine like terms

OpenStudy (anonymous):

you are right for A

OpenStudy (anonymous):

the thing about b is that x-3 doesnt equal to zero,so the answer can't be 3.but when you do the whole x^2-9 thing the answer comes out 3.so that confuses me.

OpenStudy (anonymous):

it does not come out to be 3 if you factor the numerator you get \[\frac{x^2-9}{x-3}=\frac{x+3)(x-3)}{x-3}=x+3\] so long as \(x\) is not \(3\) if \(x\) is 3, then the original expression is undefined

OpenStudy (anonymous):

oh,so then it will be 3,awesome.

OpenStudy (anonymous):

on c) i got 2x^2 -15,is it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

careful here B is NOT 3, it is \(x+3\)

OpenStudy (anonymous):

oh. and c is correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

d came as 2x^2-18.